# AdaDetectGPT: Adaptive Detection of LLM-Generated Text with Statistical Guarantees

### v1

@@ D.1 Assumptions
Assumption 3 (Stochastic dominance).
For any witness function $w\in\mathcal{W}$,
$\sum_{t}\mathbb{E}_{\widetilde{X}_{t}\sim q_{t}}w(\log q_{t}(\widetilde{X}_{t}|X_{<t}))-\mathbb{E}_{\widetilde{X}_{t}\sim p_{t}}w(\log q_{t}(\widetilde{X}_{t}|X_{<t}))\geq 0$
holds almost surely.
@@
D.2 Discussion on assumptions
Assumption 1 is commonly assumed in the the classification and inference literature; see, for instance, Audibert & Tsybakov (2007); Shi et al. (2020). For Assumption 3, the witness function is trained to maximize the mean discrepancy. When $w(x)\equiv 1$, the mean discrepancy is exactly zero. Therefore, it is reasonable to restrict the function class to those functions that yield a non-negative mean discrepancy. Assumption 4 basically requires the conditional variance of logits be asymptotically equivalent for human-authored and machine-generated passages. This assumption is not overly restrictive, as the variance discrepancy between the two types of passages is relatively small in our dataset (see Figure 2). Assumption 5 is commonly assumed in martingale central limit theorem literature, see e.g. Hall & Heyde (2014); Bolthausen (1982b). Empirical results further support the validity of Assumption 5: the sample mean of the ratio remains nearly constant as a function of $L$, while its variance (in parentheses) approaches zero as $L\to\infty$ across all three datasets (see Table S4), suggesting that this condition is practical. Furthermore, two standard hypothesis tests — the Kolmogorov–Smirnov (KS) test and the Shapiro–Wilk (SW) test — are conducted to evaluate whether the proposed statistic follows a normal distribution. As shown in Table S5, almost all $p$-values exceed 0.1 (most by a large margin), indicating that our test statistic passes normality tests in most cases. These test results also provide strong empirical support for the validity of MCLT regularity conditions.
@@ E.2 Preparatory results
See Corollary 3.1 in Hall & Heyde (2014) and Theorem 2 in Bolthausen (1982a). ∎
Lemma S1 (Convergence rate of martingale central limit theorem).
Let $\bm{X}=(X_{1},\ldots X_{n})$ be sequences of real valued random variables satisfying for all $1\leq t\leq n$
$\mathbb{E}\left\{X_{t}|X_{<t}\right\}=0\quad\textup{almost surely.}$
Let $\sigma_{t}^{2}=\mathbb{E}\left\{X_{t}^{2}\big|X_{<t}\right\}$ and $\bar{\sigma}_{t}^{2}=\mathbb{E}\left\{X_{t}^{2}\right\}$. Denote $s_{n}^{2}=\sum_{t=1}^{n}\bar{\sigma}_{t}^{2}$, $V_{n}^{2}=\sum_{t=1}^{n}\sigma_{t}^{2}/s_{n}^{2}$. Suppose $\|X_{n}\leq\gamma$ almost surely for all $n$, $s_{n}^{2}/n\to s^{2}$, and $V_{n}^{2}\to 1$ as $n\to\infty$, then
$\sup_{z\in\mathbb{R}}\left|\mathbb{P}\left(\frac{\sum_{t=1}^{n}X_{t}}{\sqrt{\sum_{t=1}^{n}\sigma_{t}^{2}}}\leq z\right)-\Phi(z)\right|\leq O\left(\frac{\log n}{\sqrt{n}s^{3}}+(\mathbb{E}|V_{n}^{2}-1|)^{1/3}\right),$
@@
From Corollary 1 in Bolthausen (1982a), we have
$\sup_{z\in\mathbb{R}}\left|\mathbb{P}\left(\frac{\sum_{t=1}^{n}X_{t}}{s_{n}}\leq z\right)-\Phi(z)\right|\leq O\left(\frac{\log n}{\sqrt{n}s^{3}}+(\mathbb{E}|V_{n}^{2}-1|)^{1/3}\right).$
@@
$\leq$ $O\left(\frac{\log n}{\sqrt{n}s^{3}}+(\mathbb{E}|V_{n}^{2}-1|)^{1/3}\right)+\sup_{z\in\mathbb{R}}|\nabla\Phi(z)|\times\mathbb{E}\left|\left(\frac{\sum_{t=1}^{n}X_{t}}{\sqrt{\sum_{t=1}^{n}\sigma_{t}^{2}}}\right)(V_{n}-1)\right|.$
Using the definition $\Phi(z)$, we know that
$\sup_{z\in\mathbb{R}}|\nabla\Phi(z)|=\sup_{z\in\mathbb{R}}\frac{1}{\sqrt{2\pi}}\exp(-z^{2}/2)\leq\frac{1}{\sqrt{2\pi}},$ (15)
and using the fact that $\|X_{t}\|_{\infty}\leq\gamma$, $s_{n}^{2}/n\to s^{2}>0$ and $V_{n}\to 1$ in probability, we obtain that
@@
$\leq$ $\left(\mathbb{E}\left\{\frac{\left(\sum_{t=1}^{n}X_{t}\right)^{2}}{s_{n}^{2}}V_{n}^{2}\right\}\mathbb{E}\left\{(V_{n}-1)^{2}\right\}\right)^{1/2}$
$=$ $O((\mathbb{E}|V_{n}^{2}-1|)^{1/2}),$
where the inequality is Cauchy inequality. Combining equations (14), (15) and (16), we obtain
$\sup_{z\in\mathbb{R}}\left|\mathbb{P}\left(\frac{\sum_{t=1}^{n}X_{t}}{\sqrt{\sum_{t=1}^{n}\sigma_{t}^{2}}}\leq z\right)-\Phi(z)\right|\leq O\left(\frac{\log n}{\sqrt{n}s^{3}}+(\mathbb{E}|V_{n}^{2}-1|)^{1/3}\right),$ (17)
@@
Suppose $X$ is a random variable. Let $\Phi(\bullet)$ be the cumulative distribution function of standard normal distribution and $\Phi^{\prime}(\bullet)$ be its derivative. Then for any random variable $X$,
@@
Since $\phi(x)=(\sqrt{2\pi})^{-1}\exp\left(-x^{2}/2\right)$, we noted that $\phi(x)\geq\phi(z_{\alpha})$ holds if and only if $z_{\alpha}\leq x\leq z_{1-\alpha}$. Therefore, if $0\leq X<z_{1-\alpha}-z_{\alpha}$, then by the mean value theorem,
@@
The following Lemma is essential to derive a sharper lower bound for Theorem 1.
Lemma S3.
Suppose $X$ is a random variable which satisfies $X\geq 0$ almost surely. Let $\Phi(\bullet)$ be the cumulative distribution function of standard normal distribution and $\phi(\bullet)$ be the density function of standard normal distribution. Then for any random variable $X$,
$\mathbb{E}\Phi(z_{\alpha}+X)\geq\sup_{\beta\in(0,\alpha]}\min\{1-\beta,\alpha+\Phi^{\prime}(z_{\beta})\mathbb{E}X\},$
Proof of Lemma S3.
Since $\Phi^{\prime}(x)=(\sqrt{2\pi})^{-1}\exp\left(-x^{2}/2\right)$, we noted that if $0<\alpha<1/2$, then for any $\beta\in(0,\alpha]$, $\Phi^{\prime}(x)\geq\Phi^{\prime}(z_{\beta})$ holds if and only if $z_{\beta}\leq x\leq z_{1-\beta}$. Therefore, if $0\leq X<z_{1-\beta}-z_{\alpha}$, then by mean value theorem,
$\Phi(z_{\alpha}+X)$ $=$ $\Phi(z_{\alpha})+\Phi^{\prime}(\xi)X\geq\alpha+\Phi^{\prime}(z_{\beta})X,$
where $\xi$ lies between $z_{\alpha}$ and $z_{1-\beta}$. Moreover, if $X\geq z_{1-\beta}-z_{\alpha}$, then $z_{\alpha}+X\geq z_{1-\beta}$, It follows that $\Phi(z_{\alpha}+X)\geq\Phi(z_{1-\beta})=1-\beta$. Therefore,
$\mathbb{E}\Phi(z_{\alpha}+X)$ $\geq$ $\mathbb{E}\min\left\{\alpha+\Phi^{\prime}(z_{\beta})X,1-\beta\right\}$
$\geq$ $\min\left\{\alpha+\Phi^{\prime}(z_{\beta})\mathbb{E}X,1-\beta\right\},$
where the last inequality follows from Jensen’s inequality. Take supremum with respect to $\beta\in(0,\alpha]$ on both sides, we obtain
$\mathbb{E}\Phi(z_{\alpha}+X)\geq\sup_{\beta\in(0,\alpha]}\min\{1-\beta,\alpha+\Phi^{\prime}(z_{\beta})\mathbb{E}X\}.$
This finishes the proof. Noted that by taking $\beta=\alpha$, the conclusion corresponds to Lemma S2. ∎
In Lemma S4 below, we provide a bound for the parameter estimation. Before doing so, we introduce some related functions
@@
where in particular we have put $\widehat{\mu}_{n}=L^{-1}\sum_{t=1}^{L}\widehat{\mu}_{t}^{(1)}-\widehat{\mu}_{t}^{(2)}$ and $\mu=L^{-1}\sum_{t=1}^{L}\mu^{(1)}_{t}-\mu^{(2)}_{t}$ with
@@
for each $t=1,\dots,L$, as well as $\widehat{\Sigma}_{n}=L^{-1}\sum_{t=1}^{L}\widehat{\Sigma}_{t}$ and $\Sigma=L^{-1}\sum_{t=1}^{L}\Sigma_{t}$ with
@@
Grant the assumptions in Section D hold. Let $\beta^{*}$ be the maximizer of the function (18) over all $\beta$’s with $\ell_{2}$ norm equal to $1$ and let $\widehat{\beta}$ be the maximizer of the empirical counterpart (19). There are absolute constants $\kappa_{1}$ and $\kappa_{2}$ depending only on the constants stated in the assumptions such that for any $z>0$ it holds that $\|\widehat{\beta}-\beta^{*}\|_{2}\leq z$ with probability at least $1-\kappa_{1}\exp\left(-\kappa_{2}d^{-5\gamma}n(\min\{z,1\})^{2}\right)$.
@@ E.3 Proof of Theorem 1
It is easy to verify that when $\bm{X}\sim p$, $T_{w}^{(1)}(\bm{X})\sigma_{q,L}/\sigma_{p,L}$ converges to standard normal distribution, using the convergence rate for martingale central limit theorem (i.e., Theorem S1), we obtain that
@@
$\qquad+O\left(\log L/\sqrt{L}\right)$
@@
$\qquad+O\left(\log L/\sqrt{L}\right).$
Take expectation on both sides, we obtain
$\mathbb{P}_{\bm{X}\sim p}\left(T_{w}(\bm{X})\leq z_{\alpha}\right)\geq\mathbb{E}\Phi(z_{\alpha}+T_{w}^{(2)}(\bm{X}))+o(1)+O(\log L/\sqrt{L}),$
because of Assumption 4. Now define $\widetilde{\sigma}_{q,L}^{2}=\mathbb{E}_{\bm{X}\sim p}\sigma_{q,L}^{2}$. Then under equal variance assumption (Assumption 4), we also have $\sigma_{q,L}-\widetilde{\sigma}_{q,L}\to 0$ in probability. It follows that for any $\epsilon>0$,
@@
$=$ $\mathbb{E}\Phi\left(z_{\alpha}+T_{w}^{(2)}(\bm{X})\frac{\sigma_{q,L}}{\widetilde{\sigma}_{q,L}-\epsilon}\right)\mathbb{I}\{|\sigma_{q,L}-\widetilde{\sigma}_{q,L}|\leq\epsilon\}$
$\geq$ $\mathbb{E}\Phi\left(z_{\alpha}+T_{w}^{(2)}(\bm{X})\frac{\sigma_{q,L}}{\widetilde{\sigma}_{q,L}-\epsilon}\right)$
$\qquad-\mathbb{E}\Phi\left((z_{\alpha}+T_{w}^{(2)}(\bm{X})\frac{\sigma_{q,L}}{\widetilde{\sigma}_{q,L}-\epsilon}\right)\mathbb{I}\{|\sigma_{q,L}-\widetilde{\sigma}_{q,L}|>\epsilon\}$
$\geq$ $\mathbb{E}\Phi\left(z_{\alpha}+T_{w}^{(2)}(\bm{X})\frac{\sigma_{q,L}}{\widetilde{\sigma}_{q,L}-\epsilon}\right)-\mathbb{P}(|\sigma_{q,L}-\widetilde{\sigma}_{q,L}|>\epsilon),$
where the first inequality is obtained due to $\Phi$ is non-negative and the last inequality holds because $T_{w}^{(2)}(\bm{X})\sigma_{q,L}>0$ under stochastic dominance assumption (Assumption 3) and $\Phi$ is bounded by $1$. Together with Lemma S2 and Assumption 4, we obtain
$\begin{split}&\mathbb{P}_{\bm{X}\sim p}\left(T_{w}(\bm{X})\leq z_{\alpha}\right)\\ \geq&\min\left\{1-\alpha,\alpha+\phi(z_{\alpha})\mathbb{E}\left\{T_{w}^{(2)}(\bm{X})\frac{\sigma_{q,L}}{\widetilde{\sigma}_{q,L}}\right\}\right\}\frac{\widetilde{\sigma}_{q,L}}{\widetilde{\sigma}_{q,L}-\epsilon}\\ &-\mathbb{P}\{|\sigma_{q,L}-\widetilde{\sigma}_{q,L}|\geq\epsilon\}+O\left(\log L/\sqrt{L}\right)+o(1).\end{split}$ (30)
Let $L\to\infty$ and using the fact that $\mathbb{E}\left\{T_{w}^{(2)}(\bm{X})\frac{\sigma_{q,L}}{\widetilde{\sigma}_{q,L}}\right\}=T_{w}^{(2*)}(\bm{X})$, we obtain that TNR is asymptotically lower bounded by $\min\{1-\alpha,\alpha+\phi(z_{\alpha})T_{w}^{(2*)}(\bm{X})\}\frac{\widetilde{\sigma}_{q,L}}{\widetilde{\sigma}_{q,L}-\epsilon}$. By taking $\epsilon\to 0$, then the conclusion of Theorem 1 follows.
@@
It is worth noting that since $T_{w}^{(2)}(\bm{X})\sigma_{q,L}>0$, a sharper lower bound can be obtained by applying Lemma S3 in the last step (inequality (30)), then follow a similar argument, we will obtain that TNR is asymptotically lower bounded by
$\sup_{0<\beta\leq\alpha}\min\{1-\beta,\alpha+\phi(z_{\beta})T_{w}^{(2*)}(\bm{X})\}.$ (31)
Noted that with any fixed $\alpha\in(0,1/2)$, the lower bound in (31) may tend to $1$ given that $T_{w}^{(2*)}(\bm{X})$ is sufficiently large, which is a sharper bound than in Theorem 1.

### v2

@@ Appendix D Technical assumptions
Conditions similar to Assumption 1 are commonly assumed; see, for instance, Audibert & Tsybakov (2007); Qian & Murphy (2011); Luedtke & Van Der Laan (2016); Shi et al. (2020a, b, 2022). Assumption 2 is mild, since the constant $\gamma$ is allowed to be arbitrarily large. Assumption 3 basically requires the conditional variance of logits be asymptotically equivalent for human-authored and machine-generated passages. This assumption is not overly restrictive, as the variance discrepancy between the two types of passages is relatively small in our dataset (see Table S5, where the ratio f the variances are very closed to 1). Assumption 4 is commonly assumed in martingale central limit theorem literature, see e.g. Bolthausen (1982); Hall & Heyde (2014). Our empirical results further support the validity of Assumption 4: the sample mean of the ratio remains nearly constant as a function of $L$, while its variance (in parentheses) approaches zero as $L\to\infty$ across all three datasets (see Table S5), suggesting that this condition is practical. Furthermore, two commonly employed hypothesis tests — the Kolmogorov–Smirnov (KS) test and the Shapiro–Wilk (SW) test — are conducted to evaluate whether the proposed statistic follows a normal distribution. As shown in Table S6, almost all $p$-values exceed 0.1 (most by a large margin), indicating that our test statistic passes normality tests in most cases. These test results also provide strong empirical support for the validity of MCLT regularity conditions.
@@ E.2 Preparatory results
We introduce three auxiliary results in this section. Theorem S6 presents a concentration inequality that is critical to establishing the learning guarantees of AdaDetectGPT in Theorem 3. Theorem S7 formally states the MCLT. Finally, Lemma S1 can be viewed as a non-asymptotic version of Theorem S7 which provides an explicit error bound on the accuracy of the normal approximation in the MCLT.
@@
See Corollary 3.1 in Hall & Heyde (2014) and Theorem 2 in Bolthausen (1982). ∎
Lemma S1 (Convergence rates in MCLT).
Let $\bm{X}=(X_{1},\ldots X_{n})$ be sequences of real valued random variables satisfying for all $1\leq t\leq n$,
$\mathbb{E}(X_{t}|X_{<t})=0\quad\textup{almost surely.}$
Let $\sigma_{t}^{2}=\mathbb{E}(X_{t}^{2}\big|X_{<t})$, $\bar{\sigma}_{t}^{2}=\mathbb{E}(X_{t}^{2})$, $s_{n}^{2}=\sum_{t=1}^{n}\bar{\sigma}_{t}^{2}$ and $V_{n}^{2}=\sum_{t=1}^{n}\sigma_{t}^{2}/s_{n}^{2}$. Suppose $|X_{n}|$ is bounded by some constant almost surely for all $n$ and $s_{n}/\sqrt{n}$ is bounded away from zero. Then
$\sup_{z\in\mathbb{R}}\left|\mathbb{P}\left(\frac{\sum_{t=1}^{n}X_{t}}{\sqrt{\sum_{t=1}^{n}\sigma_{t}^{2}}}\leq z\right)-\Phi(z)\right|=O\left(\frac{\log n}{\sqrt{n}}+(\mathbb{E}|V_{n}^{2}-1|)^{1/3}\right),$
@@
It follows from Corollary 1 of Bolthausen (1982) and the condition $s_{n}/\sqrt{n}$ is bounded away from zero that
$\sup_{z\in\mathbb{R}}\left|\mathbb{P}\left(\frac{\sum_{t=1}^{n}X_{t}}{s_{n}}\leq z\right)-\Phi(z)\right|=O\left(\frac{n\log n}{s_{n}^{3}}+(\mathbb{E}|V_{n}^{2}-1|)^{1/3}\right)=O\left(\frac{\log n}{\sqrt{n}}+(\mathbb{E}|V_{n}^{2}-1|)^{1/3}\right).$
@@
$\leq$ $O\left(\frac{\log n}{\sqrt{n}}+(\mathbb{E}|V_{n}^{2}-1|)^{1/3}\right)+\sup_{z\in\mathbb{R}}|\Phi^{\prime}(z)|\times\mathbb{E}\left|\left(\frac{\sum_{t=1}^{n}X_{t}}{\sqrt{\sum_{t=1}^{n}\sigma_{t}^{2}}}\right)(V_{n}-1)\right|.$
By the definition $\Phi(z)$, we have that
$\sup_{z\in\mathbb{R}}|\Phi^{\prime}(z)|=\sup_{z\in\mathbb{R}}\frac{1}{\sqrt{2\pi}}\exp(-z^{2}/2)\leq\frac{1}{\sqrt{2\pi}}.$ (18)
Leveraging the facts that $\sup_{t}|X_{t}|$ are upper bounded and $s_{n}/\sqrt{n}$ is lower bounded, we obtain that
@@
$\leq$ $\left(\mathbb{E}\left\{\frac{\left(\sum_{t=1}^{n}X_{t}\right)^{3/2}}{s_{n}^{3/2}}V_{n}^{3/2}\right\}\right)^{2/3}\left(\mathbb{E}\left\{(V_{n}-1)^{3}\right\}\right)^{1/3}$
$=$ $O\left(\left\{\mathbb{E}|V_{n}-1|^{2}\right\}^{1/3}\right)$
$=$ $O\left(\left\{\mathbb{E}|V_{n}^{2}-1|\right\}^{1/3}\right),$
[…]
$\sup_{z\in\mathbb{R}}\left|\mathbb{P}\left(\frac{\sum_{t=1}^{n}X_{t}}{\sqrt{\sum_{t=1}^{n}\sigma_{t}^{2}}}\leq z\right)-\Phi(z)\right|\leq O\left(\frac{\log n}{\sqrt{n}}+(\mathbb{E}|V_{n}^{2}-1|)^{1/3}\right),$ (20)
@@
Let $\Phi(\bullet)$ be the cumulative distribution function of standard normal distribution and $\Phi^{\prime}(\bullet)$ be its derivative. Then for any random variable $X$,
@@
Since $\Phi^{\prime}(x)=(\sqrt{2\pi})^{-1}\exp\left(-x^{2}/2\right)$, we noted that $\Phi^{\prime}(x)\geq\Phi^{\prime}(z_{\alpha})$ holds if and only if $z_{\alpha}\leq x\leq z_{1-\alpha}$. Therefore, if $0\leq X<z_{1-\alpha}-z_{\alpha}$, then by the mean value theorem,
@@
In Lemma S3 below, we provide an upper bound for the parameter estimation error. Before doing so, we define
@@
where $\widehat{\mu}_{n}=L^{-1}\sum_{t=1}^{L}\widehat{\mu}_{t}^{(1)}-\widehat{\mu}_{t}^{(2)}$, $\mu=L^{-1}\sum_{t=1}^{L}\mu^{(1)}_{t}-\mu^{(2)}_{t}$ and
@@
for each $t=1,\dots,L$. Similarly, define $\widehat{\Sigma}_{n}=L^{-1}\sum_{t=1}^{L}\widehat{\Sigma}_{t}$ and $\Sigma=L^{-1}\sum_{t=1}^{L}\Sigma_{t}$ where
@@
Grant the assumptions in Section D hold. Let $\beta^{*}$ be the maximizer of the function (21) over all $\beta$’s with $\ell_{2}$ norm equal to 11 and let $\widehat{\beta}$ be the maximizer of the empirical counterpart (22). There are absolute constants $\kappa_{1}$ and $\kappa_{2}$ depending only on the constants stated in the assumptions such that for any $z>0$ it holds that $\|\widehat{\beta}-\beta^{*}\|_{2}\leq z$ with probability at least $1-\kappa_{1}\exp\left(-\kappa_{2}d^{-5\gamma}n(\min\{z,1\})^{2}\right)$.
@@ E.3 Proof of Theorem 1
It is easy to verify that when $\bm{X}\sim p$, $T_{w}^{(1)}(\bm{X})\sigma_{q,L}/\sigma_{p,L}$ converges to standard normal distribution. Specifically, using Lemma S1, we obtain that
@@
$\qquad+O\left(\log L/\sqrt{L}\right)+o_{p}(1)$
@@
$\qquad+O\left(\log L/\sqrt{L}\right)+o_{p}(1),$
where the little-$o_{p}$ term arises due to the asymptotic equivalence between $\sigma_{p,L}$ and $\bar{\sigma}_{p,L}$ in Assumption 4.
Take expectation on both sides, we have by Assumption 3 that
$\mathbb{P}_{\bm{X}\sim p}\left(T_{w}(\bm{X})\leq z_{\alpha}\right)\geq\mathbb{E}\Phi(z_{\alpha}+T_{w}^{(2)}(\bm{X}))+o(1)+O(\log L/\sqrt{L}).$
Next, define $\widetilde{\sigma}_{q,L}^{2}=\mathbb{E}_{\bm{X}\sim p}\sigma_{q,L}^{2}$. It follows that $T_{w}^{(2*)}(\bm{X})=\mathbb{E}\left\{T_{w}^{(2)}(\bm{X})\frac{\sigma_{q,L}}{\widetilde{\sigma}_{q,L}}\right\}$. Under the equal variance assumption in Assumption 3, we also have $\sigma_{q,L}-\widetilde{\sigma}_{q,L}\to 0$ in probability. It follows that for any $\epsilon>0$,
@@
$\geq$ $\mathbb{E}\Phi\left(z_{\alpha}+T_{w}^{(2)}(\bm{X})\frac{\sigma_{q,L}}{\widetilde{\sigma}_{q,L}+\text{sgn}(T_{w}^{(2)})\epsilon}\right)\mathbb{I}\{|\sigma_{q,L}-\widetilde{\sigma}_{q,L}|\leq\epsilon\}$
$\geq$ $\mathbb{E}\Phi\left(z_{\alpha}+T_{w}^{(2)}(\bm{X})\frac{\sigma_{q,L}}{\widetilde{\sigma}_{q,L}+\text{sgn}(T_{w}^{(2)})\epsilon}\right)$
$\qquad-\mathbb{E}\Phi\left((z_{\alpha}+T_{w}^{(2)}(\bm{X}))\frac{\sigma_{q,L}}{\widetilde{\sigma}_{q,L}+\text{sgn}(T_{w}^{(2)})\epsilon}\right)\mathbb{I}\{|\sigma_{q,L}-\widetilde{\sigma}_{q,L}|>\epsilon\}$
$\geq$ $\mathbb{E}\Phi\left((z_{\alpha}+T_{w}^{(2)}(\bm{X}))\frac{\sigma_{q,L}}{\widetilde{\sigma}_{q,L}+\text{sgn}(T_{w}^{(2)})\epsilon}\right)-\mathbb{P}(|\sigma_{q,L}-\widetilde{\sigma}_{q,L}|>\epsilon),$
where the first inequality is obtained due to $\Phi$ is non-negative and the second inequality holds due to the monotonicity and boundedness of $\Phi$. Together with Lemma S2 and Assumption 3, we obtain
$\begin{split}&\mathbb{P}_{\bm{X}\sim p}\left(T_{w}(\bm{X})\leq z_{\alpha}\right)\\ \geq&\min\left\{1-\alpha,\alpha+\phi(z_{\alpha})\mathbb{E}\left\{T_{w}^{(2)}(\bm{X})\frac{\sigma_{q,L}}{\widetilde{\sigma}_{q,L}}\right\}\right\}\frac{\widetilde{\sigma}_{q,L}}{\widetilde{\sigma}_{q,L}+\text{sgn}(T_{w}^{(2)})\epsilon}\\ &-\mathbb{P}\{|\sigma_{q,L}-\widetilde{\sigma}_{q,L}|\geq\epsilon\}+O\left(\log L/\sqrt{L}\right)+o(1).\end{split}$ (33)
Let $L\to\infty$ and using the fact that $\mathbb{E}\left\{T_{w}^{(2)}(\bm{X})\frac{\sigma_{q,L}}{\widetilde{\sigma}_{q,L}}\right\}=T_{w}^{(2*)}(\bm{X})$, we obtain that TNR is asymptotically lower bounded by $\min\{1-\alpha,\alpha+\phi(z_{\alpha})T_{w}^{(2*)}(\bm{X})\}\frac{\widetilde{\sigma}_{q,L}}{\widetilde{\sigma}_{q,L}+\text{sgn}(T_{w}^{(2)})\epsilon}$. By taking $\epsilon\to 0$, then the conclusion of Theorem 1 follows.
@@
In fact, the equal variance condition (Assumption  3) can be relaxed. Specifically, it is not necessary for the two variance $\sigma_{q,L}^{2}$ and $\sigma_{p,L}^{2}$ to be asymptotically equivalent in probability. Rather, it suffices to require their ratio to converge to some positive constant in probability, i.e.,
$\frac{\sigma_{q,L}}{\sigma_{p,L}}\stackrel{{\scriptstyle P}}{{\rightarrow}}K_{0}.$
Since $K_{0}$ need not be 11, this proportionality condition is considerably weaker than the equal variance assumption and is more likely to hold in practice. Under this relaxed condition, following nearly identical arguments to those in Theorem 3, we can show that TNR is asymptotically lower bounded by:
$\min\left\{1-\Phi(K_{0}z_{\alpha}),\Phi(K_{0}z_{\alpha})+\phi(K_{0}z_{\alpha})K_{0}T_{w}^{(2*)}\right\}.$
This lower bound differs from the one in Theorem  3 due to the change in assumptions. However, since $\phi(K_{0}z_{\alpha})$ depends solely on $\alpha$ (not on the witness function $w$), our core conclusion – maximizing the lower bound is equivalent to maximizing $T_{w}^{(2*)}$ – remains valid. Thus, our proposed methodology remains theoretically sound.